OpenStudy (anonymous):

\dfrac{3y-8}{12} = \dfrac{y}{5} Can someone help me with this proportion?

OpenStudy (anonymous):

\[\frac{3y-8}{12} = \frac{y}{5}\]

OpenStudy (snowsurf):

When you get rational equation like this its best to get rid of the fractions first. To get rid of the fractions you want to multiply by 12 on both sides and 5 on both sides. \[12\frac{ (3y-8) }{ 12 }=\frac{ y }{ 5 }*12\] not do the same with 5.\[5(3y-8)=5\frac{ 12y }{ 5 }\]\[5(3y-8)=12y\]Use distributive law\[15y-40=12y\]Combine like terms\[3y=40\]\[y=\frac{ 40 }{ 3 }\]This fraction can be reduced to\[y=12\frac{ 4 }{ 3 }\]

OpenStudy (anonymous):

Thanks!